# The James Webb Interferometer: Space-based interferometric detections of PDS 70 b and c at 4.8 m

### v1

@@ Abstract
We observed the planet-hosting system PDS 70 with the James Webb Interferometer, JWST’s Aperture Masking Interferometric (AMI) mode within NIRISS. Observing with the F480M filter centered at 4.8 m, we simultaneously fit a geometric model to the outer disk and the two known planetary companions. We re-detect the protoplanets PDS 70 b and c at an SNR of 21 and 11, respectively. Our photometry of both PDS 70 b and c provide evidence for circumplanetary disk emission through fitting SED models to these new measurements and those found in the literature. We also newly detect emission within the disk gap at an SNR of 4, at a position angle of degrees, and an unconstrained separation within 200 mas. Follow-up observations will be needed to determine the nature of this emission. We place a 5 upper limit of mag = 7.56 on the contrast of the candidate PDS 70 d at 4.8 m, which indicates that if the previously observed emission at shorter wavelengths is due to a planet, this putative planet has a different atmospheric composition than PDS 70 b or c. Finally, we place upper limits on emission from any additional planets in the disk gap. We find an azimuthally averaged 5 upper limit of mag 7.5 at separations greater than 125 mas. These are the deepest limits to date within 250 mas at 4.8 m and the first space-based interferometric observations of this system.
@@ 1 Introduction
The planets PDS 70 b and c have been extensively imaged in the near-infrared from 1 to 3.8 m (e.g., Keppler et al., 2018; Müller et al., 2018; Mesa et al., 2019; Christiaens et al., 2019; Wang et al., 2020, 2021b; Cugno et al., 2021), with a few detections beyond 4 m to 4.8 m (Stolker et al., 2020a; Christiaens et al., 2024), as well as detections in Hα (Wagner et al., 2018; Haffert et al., 2019). Circumplanetary disk emission from PDS 70 c has been directly detected in the sub-mm with ALMA, coincident with detections in the near-infrared (Isella et al., 2019; Benisty et al., 2021). Sub-mm emission has also been observed that is seemingly associated with PDS 70 b, however it is offset from the near-infrared detections (Isella et al., 2019; Benisty et al., 2021). Additionally, emission at the L5 Lagrangian point of PDS 70 b has tentatively been detected in the sub-mm with ALMA (Balsalobre-Ruza et al., 2023). Finally, a third point source has been tentatively detected in the PDS 70 disk cavity, using VLT/SPHERE over multiple epochs (Mesa et al., 2019). The nature of the source is still not well understood as it has a distinct spectrum compared to both PDS 70 b and c, more similar to the spectrum of PDS 70 A between 1-1.6 m, implying it may be scattered light (Mesa et al., 2019). This source was tentatively re-detected at 1.9 m with JWST/NIRCam (Christiaens et al., 2024).
Many atmospheric models of massive giant planets predict a drop in flux in the mid-infrared beyond 4.5 m compared to between 2-2.5 m and 3-4 m, so observations at 4.8 m may allow for excess emission from warm circumplanetary disk (CPD) dust to be inferred. Wang et al. (2021b) showed that the VLT/NACO M′ detection of PDS 70 b from Stolker et al. (2020a) provided weak support for excess blackbody-like emission, so follow up observations at 4.8 m and beyond will allow for this tentative detection of CPD emission in the mid-infrared to be confirmed. Recently, Christiaens et al. (2024) re-detected PDS 70 b and c with the JWST/NIRCam (4.8 m) F480M filter. Their detection of PDS 70 b was in agreement with the M′ detection, but with a larger uncertainty.
In our work, we re-detect PDS 70 b and c at F480M using the James Webb Interferometer, an independent method to Christiaens et al.’s JWST/NIRCam results, and make the most precise measurement of the flux of PDS 70 b at 4.8 m, using the JWST/NIRISS F480M filter. We also detect PDS 70 c using the F480M filter at a similar precision to Christiaens et al. (2024), after including our conservative systematic noise estimation in our error budget. Wang et al. (2021b) made the first interferometric detection of PDS 70 b and c from the ground, using VLTI/GRAVITY. Here, using the power of the James Webb Interferometer, we detect PDS 70, its outer disk, and its two protoplanets, b and c. These are the first planets detected with space-based interferometry.
@@ 2 Observations and data reduction
The data were reduced using version 1.11.3 of the jwst pipeline222https://github.com/spacetelescope/jwst, from the uncal data format. We use the default calwebb_detector1 stage 1 pipeline and calwebb_image2 stage 2 pipeline. Note that we do not use the charge_migration333named undersampling_correction in the 1.11.3 version of the pipeline that we used step (Goudfrooij et al., 2024), that was designed to minimize the impact of the detector brighter-fatter effect (BFE). The charge_migration algorithm discards all data in each pixel above a set signal threshold, along with the rest of the data in the four directly adjacent neighbors. This strategy is not optimal for AMI data because, as we show in Appendix B, and has been shown in other works (e.g., Argyriou et al., 2023), charge bleeding from a bright pixel, into its neighbors, affects not only the four adjacent pixels, but also the four pixels along the diagonals, albeit somewhat less. Additionally, considering only the brightness of individual pixels is not the optimal threshold criteria because the BFE depends on the contrasts between pixels (e.g., Plazas et al., 2017). Due to the fine structure in the AMI PSF, that is somewhat undersampled at 4.8 m with the NIRISS plate scale of 66 mas, there are many parts of the PSF with large contrasts between neighboring pixels. The current pipeline BFE mitigation algorithm, therefore, is not useful for correcting AMI data. A modification to the existing pipeline, beyond the scope of this work, could be to base the threshold on the contrast between adjacent and diagonal neighboring pixels.
@@
We investigate the effect of the BFE on the data (see Appendix B for details), as it has been shown to be the limiting noise floor for AMI observations (Sivaramakrishnan et al., 2023; Kammerer et al., 2023; Ray et al., 2023; Sallum et al., 2023b). Based on our analysis, and to minimize the effect of the BFE on the calculated squared visibilities and closure phases, we discard all groups beyond where the sum of the central 33 pixels of the PSF is above 30,000 data number (DN), or 48,000 electrons, in the linearized ramp-level data. We use this cutoff, as opposed to the intensity in the central pixel, because it is more robust against changes in PSF centering. This constraint corresponds to an apparent systematic change in the calculated rate in the central pixel of 1%, and minimal bleeding into surrounding pixels (Appendix B, Figure 10). We note that this loss in signal biases our measured contrasts by significantly less than 1%, due to the central pixel accounting for 1% of the total signal in the PSF. This can be seen in Figure 1, which presents the mean PDS 70 image-plane data using only the first 28 groups. We also note that this BFE mitigation method is only possible because PDS 70 and HD 123991 are relatively faint, such that we have at least 10 groups before a peak signal level of 5800 DN. With fewer groups, the data is likely to be dominated by the 1/f noise seen in the NIRISS detector (e.g., Sallum et al., 2023b). For PDS 70 we are therefore limited to 28 out of the total 96 groups, while for HD 123991 we are limited to 10 out of the total 29 groups.
@@
The background star near PDS 70 that was originally observed by Riaud et al. (2006), and confirmed to be a background star by Hashimoto et al. (2012), is at the edge of the SUB80 array, at a separation of 2.5 arc seconds (as) from PDS 70 A. This is outside of the interferometric field-of-view, set by the shortest baseline, but the signal will still contaminate all baselines. To minimize the contribution to the calculated interferometric observables, a super-Gaussian window (), with = 30 pixels, is applied to both the PDS 70 data and the HD 123991 data.
We calculate the squared visibilities, , and closure phases, of PDS 70 and calibrator, HD 123991, using amical (Soulain et al., 2020; Soulain & Robert C. M. T., 2023). We also compute the statistical uncertainties (i.e. the standard error of the mean) of these quantities across integrations. The interferometric observables are calibrated, to remove the instrumental bias, by dividing the of PDS 70 by the of HD 123991. Similarly, the closure phases are calibrated by subtracting the closure phases of HD 123991 from those of PDS 70.
@@ 3 Methods
To model the extended disk emission, we use a simple geometrical model, as was used to succesfully reconstruct the extended emission of LkCa 15, in the near infrared, with ground based AMI (Blakely et al., 2022). A simple model is sufficient due to the sparsity of the data, and the fact that the dominant signal is expected to be from the forward scattering peak of the disk. To construct this model, we assume that all of the disk signal that we are observing is optically thick emission from the disk surface, in a similar manner to the procedure used by Stolker et al. (2016) and Sissa et al. (2018). We describe the height of the surface of the disk using a power law profile, given by
@@
The azimuthal brightness distribution is a sum of an axisymmetric component and an asymmetric component described by the Henyey-Greenstein scattering phase function (Henyey & Greenstein, 1941)
where is the scattering parameter, a free parameter constrained to be between 0 (isotropic scattering) and 1 (forward scattering), and is the azimuthal location of the scattering peak, which is fixed to the near side of the disk.
Combining Equations 2 through 6, the full disk model is given by
@@
Finally, any over-resolved emission is included in the model using a single free parameter , that contributes additional flux to the normalization of the visibilities. The full model consists of the sum of Equation 1 and the DFT of Equation 7, normalized by the total flux of the model. The complex visibilities of this 17 parameter model are given by
@@
We follow the general procedure used by Blakely et al. (2022) and Sallum et al. (2023a) of fitting geometrical models to the squared visibilities and closure phases calculated from AMI data. Our log-likelihood is the sum of an unnormalized Gaussian log-likelihood for both the closure phases and squared visibilities. For the closure phases, the uncertainties are simply the statistical uncertainties, calculated across integrations. However, for the squared visibilities, to account for a measurable drift at each baseline, observed as the signal level increases, even at low flux levels, an estimated systematic error budget of 1% is added in quadrature with the statistical uncertainties. We estimate this 1% systematic on the squared visibilities by comparing the calculated squared visibilities for PDS 70 and HD 123991 for all signal levels (we show the evolution of pixel intensities in Appendix B, Figure 10). We observe a drift from group 1 up to our cutoff signal level (groups 28 and 10 for PDS 70 and HD 123991, respectively) that is inconsistent with the calculated statistical uncertainties. We find that the difference in the evolution of the squared visibilities between PDS 70 and HD 123991 is 1% at the cutoff signal. We do not include a similar systematic for the closure phases because the evolution we see for increasing signal level is largely consistent with the calculated statistical uncertainities.
We use a Bayesian modelling approach to measure the locations and contrasts of PDS 70 b and c, as well as our model disk parameters. We estimate our model posterior using dynamic nested sampling (Higson et al., 2018), with dynesty (Speagle, 2020; Koposov et al., 2023). We use uniform priors on the contrasts of both planets, between 0 and 10 mag. We also use uniform priors on the locations of PDS 70 b and c and then we test this result against a model that uses tightly constrained Gaussian priors, henceforth referred to as GRAVITY priors, based on the planets predicted locations from Wang et al. (2021b), made using VLTI/GRAVITY astrometry. The planet location priors are given in Table 3, in Appendix A. We use priors on the the disk geometry from Keppler et al. (2018) and Keppler et al. (2019), that are described in Appendix A.
We note that there are limitations to our joint model fitting approach due to its necessarily simple nature. This means that, despite the ability of our joint-modelling approach to measure the correlations between the planet and disk parameters, the planet parameter uncertainties derived from our nested sampling fitting do not capture the full uncertainties in the physical parameters that we are trying to measure. To provide conservative estimates on the uncertainties of our derived parameters for PDS 70 b and c, we estimate a systematic uncertainty on their separations, position angles and contrasts. We use an injection/recovery approach similar to what is used by Stolker et al. (2024). Because of the non-linearity of closure phases (especially in the moderate contrast regime, where the visibility amplitudes are not 1), we mimic the injection of point sources into the data by injecting a companion with a negative flux into the model complex visibilities before we calculate the model closure phase and squared visibilities. We inject the companion at the median separation and (negative) contrast of PDS 70 b and c found from our nested sampling fits at position angles between 0 and 360 degrees, in increments of 5 degrees. We recover the parameters using nested sampling for each case, while fixing the disk parameters to the median parameters from the original nested sampling fit. To estimate the systematic uncertainty on the planet parameters, we use the root-mean-square deviation of the median estimated parameters from the true parameters, at each position angle.
The mean parameters for the locations and contrasts of PDS 70 b and c, along with the standard deviation of the distribution calculated with nested sampling with the estimated systematic uncertainty added in quadrature are shown in Table 1. For PDS 70 b, we estimate a systematic in the contrast of 0.08 mags, 3.4 mas in separation and 1.2 degrees in position angle. Only the contrast systematic has a significant contribution to our total error budget, compared to the uncertainties estimated with nested sampling (Appendix A, Table 3). For PDS 70 c, we find a contrast systematic of 0.08 mags, a separation systematic of 7.0 mas, and a position angle systematic of 0.8 degrees. Both the contrast and separation systematic uncertainties have a non-negligible contribution to the total error budget, indicating that there is some level of degeneracy between the disk signal and a companion with the contrast that we estimate for PDS 70 c that we are able to capture using this injection/recovery procedure.
To convert the measured contrasts to fluxes, we model the spectral energy distribution of the unresolved stellar plus inner disk emission using a power law, assuming that any spectral features over the range of the F480M filter (4.6-5.0 m) are negligible, using the same approach as is done by Stolker et al. (2020a). We fit the power law, in log-log space, to 2MASS (Skrutskie et al., 2006) J, H and Ks as well as WISE (Wright et al., 2010) W1 and W2 magnitudes that we convert to fluxes using species (Stolker et al., 2020b). We employ nested sampling (Skilling, 2004), using dynesty, to estimate the posterior of the power-law model. To account for the outer disk flux, which we measure in our fits to be 4% of the stellar flux, we scale the model spectrum by 0.96, finding a stellar magnitude of 7.56 0.02, in F480M. To convert the measured contrasts of the planets to physical units we integrate the model spectrum across the F480M filter transmission profile. To estimate the uncertainties on the derived fluxes we do this for all combinations of the planet contrast posterior samples and 500 randomly drawn samples from the power law model posterior, and calculate confidence intervals from this distribution.
| Separation | Position Angle | Contrast | Flux | SNR | |
| (mas) | (∘) | (mag) | (W m-2 m-1) | ||
| PDS 70 b | 21.1 | ||||
| PDS 70 c | 10.6 |
Notes: We report the mean and the standard deviation of the marginalized posteriors, with the systematic uncertainties estimated from the injection/recovery added in quadrature with the nested sampling uncertainties. The SNR is defined here as the mean of the contrast posterior samples divided by its standard deviation from nested sampling only (displayed in Appendix A, Table 3).

### v2

@@ Abstract
We observed the planet-hosting system PDS 70 with the James Webb Interferometer, JWST’s Aperture Masking Interferometric (AMI) mode within NIRISS. Observing with the F480M filter centered at 4.8 m, we simultaneously fit geometrical models to the outer disk and the two known planetary companions. We re-detect the protoplanets PDS 70 b and c at an SNR of 14.7 and 7.0, respectively. Our photometry of both PDS 70 b and c provides tentative evidence of mid-IR circumplanetary disk emission through fitting SED models to these new measurements and those found in the literature. We also newly detect emission within the disk gap at an SNR of 4, at a position angle of degrees, and an unconstrained separation within 200 mas. Follow-up observations will be needed to determine the nature of this emission. We place a 5 upper limit of 208 10 Jy on the flux of the candidate PDS 70 d at 4.8 m, which indicates that if the previously observed emission at shorter wavelengths is due to a planet, this putative planet has a different atmospheric composition than PDS 70 b or c. Finally, we place upper limits on emission from any additional planets in the disk gap. We find an azimuthally averaged 5 contrast upper limit 7 magnitudes at separations greater than 110 mas. These are the deepest limits to date within 250 mas at 4.8 m and the first space-based interferometric observations of this system.
@@ 1 Introduction
The planets PDS 70 b and c have been extensively imaged in the near-infrared from 1 to 3.8 m (e.g., Keppler et al., 2018; Müller et al., 2018; Mesa et al., 2019; Christiaens et al., 2019; Wang et al., 2020, 2021b; Cugno et al., 2021), with a few detections beyond 4 m to 4.8 m (Stolker et al., 2020; Christiaens et al., 2024), as well as detections in Hα (Wagner et al., 2018; Haffert et al., 2019). Circumplanetary disk emission from PDS 70 c has been directly detected in the sub-mm with ALMA, coincident with detections in the near-infrared (Isella et al., 2019; Benisty et al., 2021). Sub-mm emission has also been observed that is seemingly associated with PDS 70 b, however it is offset from the near-infrared detections (Isella et al., 2019; Benisty et al., 2021). Additionally, emission at the L5 Lagrangian point of PDS 70 b has tentatively been detected in the sub-mm with ALMA (Balsalobre-Ruza et al., 2023). Finally, a third point source has been tentatively detected in the PDS 70 disk cavity, using VLT/SPHERE over multiple epochs (Mesa et al., 2019). The nature of the source is still not well understood as it has a distinct spectrum compared to both PDS 70 b and c, more similar to the spectrum of PDS 70 A between 1-1.6 m, implying it may be scattered light (Mesa et al., 2019). This source was tentatively re-detected at 1.9 m with JWST/NIRCam (Christiaens et al., 2024).
Beyond 4 m, emission from warm (100 K) circumplanetary disk (CPD) material is expected to contribute comparably (and greater at longer wavelengths) to atmospheric emission (e.g., Wang et al., 2021b; Cugno et al., 2024). So, observations into the mid infrared may allow for the presence of circumplanetary disk emission to be inferred by looking for excess emission compared to atmospheric models. Wang et al. (2021b) showed that the VLT/NACO M′ detection of PDS 70 b from Stolker et al. (2020) provided weak support for excess blackbody-like emission, so follow up observations at 4.8 m and beyond will allow for this tentative detection of CPD emission in the mid-infrared to be confirmed. Recently, Christiaens et al. (2024) re-detected PDS 70 b and c with the JWST/NIRCam (4.8 m) F480M filter. Their detection of PDS 70 b was in agreement with the M′ detection, but with a larger uncertainty.
In our work, we re-detect PDS 70 b and c at F480M using the Aperture Masking Interferometry mode on JWST/NIRISS, an independent method to the JWST/NIRCam results, presented by Christiaens et al. (2024), and make the most precise measurement of the flux of PDS 70 b at 4.8 m, using the JWST/NIRISS F480M filter. We also detect PDS 70 c using the F480M filter at a similar precision to Christiaens et al. (2024). Here, using the power of NIRISS/AMI, as introduced in these previous works using AMI: Lau et al. (2024); Ray et al. (2023); Sallum et al. (2024), we detect PDS 70, its outer disk, and its two protoplanets, b and c. These are the first planets detected with space-based interferometry.
@@ 2 Observations and data reduction
The data were reduced using version 1.11.3 of the jwst pipeline222https://github.com/spacetelescope/jwst (Bushouse et al., 2023), with Calibration Reference Data System context jwst_1110.pmap, from the uncal data format. We use the default calwebb_detector1 stage 1 pipeline and calwebb_image2 stage 2 pipeline, skipping only the IPC, photom and resample steps. Note that we do not use the charge_migration333named undersampling_correction in the 1.11.3 version of the pipeline that we used step (Goudfrooij et al., 2024), that was designed to minimize the impact of the detector brighter-fatter effect (BFE). The charge_migration algorithm discards all data in each pixel above a set signal threshold, along with the rest of the data in the four directly adjacent neighbors. This strategy is not optimal for AMI data because, as we show in Appendix B, and has been shown in other works (e.g., Argyriou et al., 2023), charge bleeding from a bright pixel, into its neighbors, affects not only the four adjacent pixels, but also the four pixels along the diagonals, albeit somewhat less. Additionally, considering only the brightness of individual pixels is not the optimal threshold criteria because the BFE depends on the contrasts between pixels (e.g., Plazas et al., 2017). Due to the fine structure in the AMI PSF, that is somewhat undersampled at 4.8 m with the NIRISS plate scale of 66 mas, there are many parts of the PSF with large contrasts between neighboring pixels. The current pipeline BFE mitigation algorithm, therefore, is not useful for correcting AMI data. A modification to the existing pipeline, beyond the scope of this work, could be to base the threshold on the contrast between adjacent and diagonal neighboring pixels.
@@
We investigate the effect of the BFE on the data (see Appendix B for details), as it has been shown to be the limiting noise floor for AMI observations (Sivaramakrishnan et al., 2023; Kammerer et al., 2023; Ray et al., 2023; Sallum et al., 2024). Based on our analysis, and to minimize the effect of the BFE on the calculated squared visibilities and closure phases, we discard all groups beyond where the sum of the central 33 pixels of the PSF is above 30,000 data number (DN), or 48,000 electrons, in the linearized ramp-level data. We use this cutoff, as opposed to the intensity in the central pixel, because it is more robust against changes in PSF centering. This constraint corresponds to an apparent systematic change in the calculated rate in the central pixel of 1%, and minimal bleeding into surrounding pixels (Appendix B, Figure 11). We note that this loss in signal biases our measured contrasts by significantly less than 1%, due to the central pixel accounting for 1% of the total signal in the PSF. This can be seen in Figure 1, which presents the mean PDS 70 image-plane data using only the first 28 groups. We also note that this BFE mitigation method is only possible because PDS 70 and HD 123991 are relatively faint, such that we have at least 10 groups before a peak signal level of 5800 DN. With fewer groups, the data is likely to be dominated by the 1/f noise seen in the NIRISS detector (e.g., Sallum et al., 2024). For PDS 70 we are therefore limited to 28 out of the total 96 groups, while for HD 123991 we are limited to 10 out of the total 29 groups.
@@
The background star near PDS 70 that was originally observed by Riaud et al. (2006), and confirmed to be a background star by Hashimoto et al. (2012), is at the edge of the SUB80 array, at a separation of 2 arc seconds (as) from PDS 70 A. This is outside of the interferometric field-of-view, set by the shortest baseline, but the signal will still contaminate all baselines. To minimize the contribution to the calculated interferometric observables, a super-Gaussian window (), with = 30 pixels, is applied to both the PDS 70 data and the HD 123991 data.
We calculate the squared visibilities, , and closure phases, of PDS 70 and calibrator, HD 123991, using amical (Soulain et al., 2020; Soulain & Robert C. M. T., 2023). We also compute the statistical uncertainties (i.e. the standard error of the mean) of these quantities across integrations. The interferometric observables are calibrated, to remove the instrumental bias, by dividing the of PDS 70 by the of HD 123991. Similarly, the closure phases are calibrated by subtracting the closure phases of HD 123991 from those of PDS 70. The calibrated closure phases and squared visibilities are shown in Figure 2.
@@ 3 Methods
To model the extended disk emission, we use a simple geometrical model, as has been used to successfully reconstruct the extended emission of LkCa 15, in the near infrared, with ground based AMI (Blakely et al., 2022; Sallum et al., 2023). A simple model is sufficient due to the sparsity of the data, and the fact that the dominant signal is expected to be from the forward scattering peak of the disk. To construct this model, we assume that all of the disk signal that we are observing is either scattered light or optically thick emission from the disk surface, in a similar manner to the procedure used by Stolker et al. (2016) and Sissa et al. (2018). We describe the height of the surface of the disk using a power law profile, given by
@@
The azimuthal brightness distribution is the sum of an axisymmetric component and an asymmetric component. For the asymmetric component, we explore three independent parametrizations: a Henyey-Greenstein model (Henyey & Greenstein, 1941), a Gaussian model, and a power law model. The Henyey-Greenstein model is given by
where is the scattering parameter, a free parameter constrained to be between 0 (isotropic scattering) and 1 (forward scattering), is the azimuthal angle in the disk mid-plane (defined by the coordinates x and y), and is the azimuthal location of the scattering peak, which is fixed to the minor axis on the near side of the disk. The Gaussian model is given by
where describes the azimuthal extent of the disk emission. The power law model is given by
where describes the azimuthal extent of the disk emission.
Combining Equations 2 through 5, and any one of scattering phase functions given by Equations 6, 7, or 8, the full disk model is given by
@@
Finally, any over-resolved emission is included in the model using a single free parameter , that contributes additional flux to the normalization of the visibilities. The full model consists of the sum of Equation 1 and the DFT of Equation 9, normalized by the total flux of the model. The complex visibilities of these 17 parameter models are given by
@@
To fit our model to the data, we further construct a more robust set of observables from the squared visibilities and closure phases, following a similar approach to the Bayesian analysis in Xuan et al. (2024) (and described in more detail in Thompson et al. (in prep.)). We calculate self-calibrating log-closure amplitudes, which calibrate out hole-based gain terms affecting the squared visibilities. We calculate the log-closure amplitude uncertainties using a Monte Carlo approach, by calculating the standard deviation of 1,000,000 noise realizations from the statistical uncertainties of the squared visibilities. We also consider correlations between shared baselines, using analytic correlation matrices for the log-closure amplitudes and closure phases. For the closure phases, we use the closure phase correlation matrix, , described by Kammerer et al. (2020) (with no spectral component). We construct a similar log-closure amplitude correlation matrix, , following Blackburn et al. (2020). The closure phase and log-closure amplitude correlation matrices can be constructed using Equations B5 and B14 derived by Blackburn et al. (2020), respectively, setting for closure phases and for log-closure amplitudes. Finally, we construct linearly independent closure phase and log-closure amplitude observables. We calculate projection matrices by taking the Singular Value Decomposition (SVD) of , where is the design matrix of either the closure phases or log-closure amplitudes, as defined by Blackburn et al. (2020), and only taking the non-zero singular values. This is similar to what is described by Ireland (2013) for closure phases. We henceforth refer to these new observables as kernel phases and log kernel amplitudes. The advantage of using the “kernel” observables is twofold. By using closure phases, we are “double counting” the data in the likelihood calculation, leading to overly confident results and underestimated parameter uncertainties. The kernel approach also allows for the covariance matrices, described above, to be used in the likelihood, as the matrices are rank-deficient, whereas this is not the case in the new basis.
We use a Bayesian modelling approach to measure the locations and contrasts of PDS 70 b and c, as well as our disk model parameters. We estimate our model posterior using dynamic nested sampling (Higson et al., 2018), with dynesty (Speagle, 2020; Koposov et al., 2023). We use uniform priors on the contrasts, separations and position angles of both planets. Our contrast prior ranges from 0 to 10 magnitudes. Our location priors for PDS 70 b are from 100-250 mas in separation and 70-160 degrees in position angle. Similarly, for the position of PDS 70 c, we use a range of 100-250 mas in separation and 230-320 degrees in position angle. We use priors on the the disk geometry from Keppler et al. (2018) and Keppler et al. (2019), that are described in Appendix A. To account for an unknown level of miscalibration seen in the squared visibilities, which is also present in the log-closure amplitudes, we include a systematic uncertainty, , term in our model, added in quadrature with the statistical uncertainty, that is recovered along with our star plus disk plus planet model parameters.
We estimate parameters for both PDS 70 b and c using Bayesian model averaging over the three asymmetric scattering models we explore. We follow the approach used by Nixon et al. (2024) and Nasedkin et al. (2024), of assuming that each model is equal likely apriori and thus combine the posterior distributions by weighting each model by their Bayesian evidences (listed in Table 2 in Appendix A).
To convert the measured contrasts to flux measurements, we use aperture photometry. We follow the procedure outlined in Section 2, of reducing and cleaning the data. The only difference being that we now apply the photom step, so that the data is in units of surface brightness. We crop the data to a size of 6767 pixels, with PDS 70 A at the center. We calculate the total flux within a 67 pixel diameter circular aperture and recover a total flux of 128.6 mJy. We note that this value includes resolved disk flux, which can be clearly seen by looking at the squared visibilities decrement from 1, seen at all baselines, in Figure 2. From the visibility amplitudes, we calculate a resolved flux component of 5.2 0.6 %, independent of baseline. Additionally, our photometry is missing a non-negligible amount of flux that is outside of the field of view. We calculate a correction factor by calculating PSF models using WebbPSF (Perrin et al., 2012). We calculate a 6767 pixel model and 20002000 pixel model, both with the native NIRISS pixel size. We calculate the correction factor by dividing the total flux within a 2000 pixel diameter circular aperture for the full PSF and a 67 pixel diameter for the cutoff PSF, giving a correction factor of 1.12. After applying these correction factors, we find a flux of PDS 70 A (and the unresolved inner disk emission) of 136.4 6.8 mJy, assuming a 5% absolute flux calibration uncertainty444https://jwst-docs.stsci.edu/jwst-calibration-status/niriss-calibration-status/niriss-imaging-calibration-status.